# Nonparametric Smoothing of Directional and Axial Data

### v1

@@ Abstract.
Bingham distributions, exponential family, Galileo mission, generalized linear model, local polynomial modelling, stereographic projection, von Mises–Fisher distributions.
@@ 1 Introduction
The starting point for the present paper are data sets from Planetary Sciences. The Galileo Mission yielded raw imaging data from the jovian111of the planet Jupiter moons Ganymede and Europa. Both moons have a global subsurface ocean with an ice shell on top. From pictures of linear surface features on the icy surface, the physicists extracted observation pairs $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$, $1\leq i\leq n$, consisting of a center position $\boldsymbol{X}_{\!i}$ on the surface at which a feature, presumably a crack in the ice layer, could be identified, and the direction $\boldsymbol{V}_{\!\!i}$ of that crack. Identifying the surface of the moon with the unit sphere $\mathbb{S}^{2}$ of $\mathbb{R}^{3}$, we are talking about points $\boldsymbol{X}_{\!i}\in\mathbb{S}^{2}$ and $\boldsymbol{V}_{\!\!i}\in\mathbb{S}^{2}\cap\boldsymbol{X}_{\!i}^{\perp}$, where $\boldsymbol{V}_{\!\!i}$ represents the axis $\mathbb{R}\boldsymbol{V}_{\!\!i}$. Figures 1 and 2 illustrate this process for the moon Europa. Figure 1 shows a superposition of a basemap with a particular region highlighted. In addition one sees some fine structures on the icy surface which become visible when zooming in. By means of tailored classic and deep learning algorithms (Haslebacher et al., 2024, 2025) one can extract data pairs $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$ as described before, see Figure 2. The whole data set contains $5623$ observation pairs in $19$ different regions, but we only show $2184$ of these pairs. The region in the foreground (black dots, green lines) corresponds to the region with red boundary in Figure 1. It contains $2656$ observations $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$, and we show a subset of $200$ pairs.
The task was to smooth these noisy data and to determine for various points $\boldsymbol{x}_{o}\in\mathbb{S}^{2}$ whether there was a preferred axis direction of ice cracks close to $\boldsymbol{x}_{o}$. The existence and strength of preferred directions allows conclusions about physical properties of the moon, e.g. the link of tidal forces to ice cracks (Rhoden and Hurford, 2013). Translated into a mathematical task, we wanted to determine for such a point $\boldsymbol{x}_{o}$ a Bingham distribution (Bingham, 1974) on the unit circle $\mathbb{S}^{2}\cap\boldsymbol{x}_{o}^{\perp}$ of the tangent space $\boldsymbol{x}_{o}^{\perp}$ of the unit sphere at $\boldsymbol{x}_{o}$.
@@
Two general references about directional data, including regression methods, are the monograph of Mardia and Jupp (2000) and the review paper of Pewsey and García-Portugués (2021). These references describe and cite numerous approaches to regression with directional response. Some of these approaches involve transformations of the directional response into points in some Euclidean space, but the choice of this transformation is somewhat ad hoc. Let us restrict our attention to models in which $\boldsymbol{Y}_{\!\!i}\in\mathbb{S}^{d-1}$ has a von Mises–Fisher distribution conditional on $\boldsymbol{X}_{\!i}$. Many authors characterize a von Mises–Fisher distribution via a direction parameter $\boldsymbol{\beta}\in\mathbb{S}^{d-1}$ and a concentration parameter $\kappa\geq 0$ which is often viewed as a nuisance parameter, whereas $\boldsymbol{\beta}$ is the parameter of interest. We propose to replace such a parameter pair $(\boldsymbol{\beta},\kappa)$ with the vector $\boldsymbol{z}=\kappa\boldsymbol{\beta}\in\mathbb{R}^{d}$. Since $\kappa=\|\boldsymbol{z}\|$ and $\boldsymbol{\beta}=\kappa^{-1}\boldsymbol{\beta}$, this description is equivalent, and dealing with parameters in a Euclidean space is more convenient, particularly in generalized linear models.
The remainder of this paper is structured as follows. Section 2 starts with a short recap of von Mises–Fisher (vMF) distributions and Bingham (Bh) distributions on the unit sphere $\mathbb{S}^{d-1}$ of $\mathbb{R}^{d}$. Then we describe a very useful connection between these two families for dimension $d=2$. Finally, in view of the regression methods derived later, we provide some details about the numerical computation of the mean vector and covariance matrix of a von Mises–Fisher distribution. This involves exact formulae as well as approximations for parameter vectors with large norm and is potentially of independent interest.
Section 3 describes regression methods for data with directional response vectors. As a parametric approach, we propose generalized linear models (GLMs). For a thorough introduction and overview of this topic we refer to McCullagh and Nelder (1989). Thereafter we describe a modification of this parametric approach to local parametric modelling in the spirit of Fan et al. (1998). The latter method is illustrated with simulated data.
In Section 4 we explain how to apply the nonparametric methods of Section 3 to the specific data mentioned in the beginning, a key tool being stereographic projections of the covariate vectors $\boldsymbol{X}_{\!i}\in\mathbb{S}^{2}$ onto $\mathbb{R}^{2}$ and suitable transformations of the accompanying axes $\boldsymbol{V}_{\!\!i}\in\mathbb{S}^{2}\cap\boldsymbol{X}_{\!i}^{\perp}$ into points on the unit circle $\mathbb{S}^{1}$. The methods are illustrated with specific data from the jovian moon Europa.
Section 5 contains further technical details and proofs.
@@ 2.1 Von Mises–Fisher distributions
The family $(\mathrm{vMF}(\boldsymbol{z}))_{\boldsymbol{z}\in\mathbb{R}^{d}}$ is a natural exponential family, and it is well-known that for a random vector $\boldsymbol{Y}\sim\mathrm{vMF}(\boldsymbol{z})$, its mean vector and covariance matrix are given by the gradient and Hessian matrix of $\gamma$,
@@
Other well-known facts are that the mapping $\boldsymbol{\mu}:\mathbb{R}^{d}\to\{\boldsymbol{x}\in\mathbb{R}^{d}:\|\boldsymbol{x}\|<1\}$ is a diffeomorphism, and by means of a Lagrange argument one can show that the density $f_{\boldsymbol{z}}$ maximizes differential Shannon entropy
@@ 2.3 The special case of $d=2$
If we identify any vector in $\mathbb{R}^{2}$ with a complex number, then $\boldsymbol{y}(\boldsymbol{v})=\boldsymbol{v}^{2}$. In particular, this shows immediately that $\|\boldsymbol{y}(\boldsymbol{v})\|=1$. Hence there is a one-to-one correspondence between axes in $\mathbb{R}^{2}$ and the unit circle $\mathbb{S}^{1}$, and specifying the expectation $\int\boldsymbol{v}\boldsymbol{v}^{\top}\,Q(\mathrm{d}\boldsymbol{v})$ for a distribution $Q$ on $\mathbb{S}^{1}$ is equivalent to specifying the expectation $\int\boldsymbol{y}(\boldsymbol{v})\,Q(\mathrm{d}\boldsymbol{v})$. Note further that any symmetric matrix $\boldsymbol{W}\in\mathbb{R}^{2\times 2}$ may be written as
$\boldsymbol{W}\ =\ \begin{bmatrix}z_{1}&z_{2}\\ z_{2}&-z_{1}\end{bmatrix}$
@@
Furthermore, if $\boldsymbol{V}\sim M$, then $\boldsymbol{y}(\boldsymbol{V})\sim M$ too, and these findings imply the very useful fact that
$\boldsymbol{V}\sim\mathrm{Bh}(\boldsymbol{W})\quad\text{if and only if}\quad\boldsymbol{y}(\boldsymbol{V})\ \sim\ \mathrm{vMF}(\boldsymbol{z}).$ (2)
@@
$\tilde{f}_{\boldsymbol{w}}(\theta)\ :=\ e^{\kappa\cos(2(\theta-\beta))-\tilde{\gamma}(\kappa)}$ (3)
@@
$\tilde{\gamma}(\kappa)\ :=\ \log\int e^{\kappa y_{1}}\,M(\mathrm{d}\boldsymbol{y})\ =\ \log\Bigl{(}\frac{1}{2\pi}\int_{0}^{2\pi}e^{\kappa\cos(2(\theta-\beta))}\,\mathrm{d}\theta\Bigr{)}.$
@@
$\boldsymbol{w}\ :=\ \kappa[\cos(\beta),\sin(\beta)]^{\top}$
@@
Figure 3 depicts the distribution $\mathrm{Bh}(\boldsymbol{w})$ for two different vectors $\boldsymbol{w}=\kappa[\cos(\beta),\sin(\beta)]^{\top}$. One possible representation is the corresponding angular density $\tilde{f}$ in (3). Another possibility is to draw an “axial histogram”, that is, the region surrounded by the curve $\theta\mapsto\tilde{f}(\theta)^{1/2}[\cos(\theta),\sin(\theta)]^{\top}$, leading to a symmetric but non-circular “pie” (unless $\kappa=0$). The area of any “slice” of this pie with boundary angles $0\leq\theta_{1}<\theta_{2}\leq 2\pi$ equals $2\pi$ times the probability that $\tilde{V}\in[\theta_{1},\theta_{2})$.
@@
2.4 Computational aspects of vMF distributions
The function $\gamma$.
$h_{d}(u)\ :=\ C_{d}(1-u^{2})^{a_{d}-1},\quad u\in(-1,1),$ (4)
where $a_{d}:=(d-1)/2$ and $C_{d}:=1/B(1/2,a_{d})$ with the beta function $B(\cdot,\cdot)$. This implies that
$\gamma(\boldsymbol{z})\ =\ \tilde{\gamma}_{d}(\|\boldsymbol{z}\|),$
where
Mean and Covariance of $\mathrm{vMF}(\boldsymbol{z})$.
Note first that symmetry considerations reveal that $\boldsymbol{Y}_{\!\!\boldsymbol{0}}\sim\mathrm{vMF}(\boldsymbol{0})$ satisfies
$\boldsymbol{\mu}(\boldsymbol{z})$ $=\ \operatorname{\mathbb{E}}(U_{t})\boldsymbol{v},$
$\boldsymbol{\Sigma}(\boldsymbol{z})$ $=\ \mathrm{Var}(U_{t})\,\boldsymbol{v}\boldsymbol{v}^{\top}+\frac{1-\operatorname{\mathbb{E}}(U_{t}^{2})}{d-1}(\boldsymbol{I}_{d}-\boldsymbol{v}\boldsymbol{v}^{\top}).$
Note that
with $G_{d}^{(\ell)}$ denoting the $\ell$-th derivative of $G_{d}$. Thus,
From the series expansion of $G_{d}(t)$ and the representations $\operatorname{\mathbb{E}}(U_{t})=\tilde{\gamma}_{d}^{\prime}(t)$, $\mathrm{Var}(U_{t})=\tilde{\gamma}_{d}^{\prime\prime}(t)$ one can deduce expansions of $\gamma(\boldsymbol{z})$, $\boldsymbol{\mu}(\boldsymbol{z})$ and $\boldsymbol{\Sigma}(\boldsymbol{z})$ for $\boldsymbol{z}$ close to $\boldsymbol{0}$.
Lemma 1.
As $\boldsymbol{z}\to 0$,
$\gamma(\boldsymbol{z})$ $=\ \frac{\|\boldsymbol{z}\|^{2}}{2d}-\frac{\|\boldsymbol{z}\|^{4}}{4d(d+2)}+O(\|\boldsymbol{z}\|^{6}),$
$\boldsymbol{\mu}(\boldsymbol{z})$ $=\ \frac{1}{d}\Bigl{(}1-\frac{\|\boldsymbol{z}\|^{2}}{d(d+2)}\Bigr{)}\boldsymbol{z}+O(\|\boldsymbol{z}\|^{5}),$
$\boldsymbol{\Sigma}(\boldsymbol{z})$ $=\ \frac{1}{d}\Bigl{(}1-\frac{\|\boldsymbol{z}\|^{2}}{d(d+2)}\Bigr{)}\,\boldsymbol{I}_{d}-\frac{2}{d^{2}(d+2)}\boldsymbol{z}\boldsymbol{z}^{\top}+O(\|\boldsymbol{z}\|^{4}).$
Numerical computation of $G_{d}^{(\ell)}(t)$ for moderate values of $t$.
Note first that
Consequently,
$G_{d}^{\prime}(t)\ =\ \frac{t}{d}\,G_{d+2}(t).$ (6)
From this formula, one can proceed inductively. In particular,
$G_{d}^{\prime\prime}(t)\ =\ \frac{1}{d}\,G_{d+2}(t)+\frac{t^{2}}{d(d+2)}\,G_{d+4}(t).$
For large values of $t$, computing $G_{d}(t)$, $G_{d+2}(t)$ and $G_{d+4}(t)$ via their series expansion becomes problematic. Here we resort to approximation formulae. The following result is proved in Section 5.
Lemma 2.
@@ 3 Regression methods for directional data
3.1 Generalized linear models
@@
Specifically, suppose that each function $\boldsymbol{f}\in\mathcal{F}$ is equal to $\boldsymbol{f}=(f_{k})_{k=1}^{d}$ with all components $f_{k}$ belonging to the same finite-dimensional linear space $\mathcal{F}^{o}$ of functions $f^{o}:\mathcal{X}\to\mathbb{R}$. If $f^{o}_{1},\ldots,f^{o}_{r}$ is a basis of $\mathcal{F}^{o}$, then any function $\boldsymbol{f}\in\mathcal{F}$ can be represented as
$L(\boldsymbol{\Theta})\ :=\ \ell(\boldsymbol{\Theta}\boldsymbol{F}^{o})\ =\ \sum_{i=1}^{n}\bigl{(}\gamma(\boldsymbol{\Theta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i}))-\boldsymbol{Y}_{\!\!i}^{\top}\boldsymbol{\Theta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})\bigr{)}.$
Since $\boldsymbol{\mu}(\boldsymbol{z})$ and $\boldsymbol{\Sigma}(\boldsymbol{z})$ are the gradient and Hessian matrix of $\gamma$ at $\boldsymbol{z}$, for $\boldsymbol{\Theta},\boldsymbol{\Delta}\in\mathbb{R}^{d\times r}$,
$L(\boldsymbol{\Theta}+\boldsymbol{\Delta})\ =$ $L(\boldsymbol{\Theta})+\sum_{i=1}^{n}(\boldsymbol{\mu}(\boldsymbol{\Theta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i}))-\boldsymbol{Y}_{\!\!i})^{\top}\boldsymbol{\Delta}\boldsymbol{X}_{\!i}$
$+\ \frac{1}{2}\sum_{i=1}^{n}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})^{\top}\boldsymbol{\Delta}^{\top}\boldsymbol{\Sigma}(\boldsymbol{\Theta}\boldsymbol{X}_{\!i})\boldsymbol{\Delta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})+O(\|\boldsymbol{\Delta}\|_{F}^{3})$
$\sum_{i=1}^{n}(\boldsymbol{\mu}(\boldsymbol{\Theta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i}))-\boldsymbol{Y}_{\!\!i})^{\top}\boldsymbol{\Delta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})\ =\ \langle\boldsymbol{\Delta},\boldsymbol{G}(\boldsymbol{\Theta})\rangle$
with the gradient matrix
$\boldsymbol{G}(\boldsymbol{\Theta})\ :=\ \sum_{i=1}^{n}(\boldsymbol{\mu}(\boldsymbol{\Theta}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i}))-\boldsymbol{Y}_{\!\!i})\boldsymbol{X}_{\!i}^{\top}\ \in\ \mathbb{R}^{d\times r}.$ (7)
where
$\sum_{i=1}^{n}\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})\boldsymbol{F}^{o}(\boldsymbol{X}_{\!i})^{\top}\otimes\boldsymbol{\Sigma}(\boldsymbol{\Theta}\boldsymbol{X}_{\!i}).$ (8)

### v4

@@ Abstract.
Bingham distributions, exponential family, Galileo space mission, generalized linear model, local polynomial modelling, stereographic projection, von Mises–Fisher distributions.
@@ 1 Introduction
The starting point for the present paper are data sets from Planetary Sciences. The Galileo Mission yielded raw imaging data from the moons Ganymede and Europa of the planet Jupiter. Both moons have a global subsurface ocean with an ice shell on top. From pictures of linear surface features on the icy surface, the physicists extracted observation pairs $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$, $1\leq i\leq n$, consisting of a center position $\boldsymbol{X}_{\!i}$ on the surface at which a feature, presumably a crack in the ice layer, could be identified, and the direction $\boldsymbol{V}_{\!\!i}$ of that crack. Identifying the surface of the moon with the unit sphere $\mathbb{S}^{2}$ of $\mathbb{R}^{3}$, we are talking about points $\boldsymbol{X}_{\!i}\in\mathbb{S}^{2}$ and $\boldsymbol{V}_{\!\!i}\in\mathbb{S}^{2}\cap\boldsymbol{X}_{\!i}^{\perp}$, where $\boldsymbol{V}_{\!\!i}$ represents the axis $\mathbb{R}\boldsymbol{V}_{\!\!i}$. Figures 1 and 2 illustrate this process for the moon Europa. Figure 1 shows a superposition of a basemap with a particular region highlighted. In addition one sees some fine structures on the icy surface which become visible when zooming in. By means of tailored classic and deep learning algorithms (Haslebacher et al., 2024, 2025) one can extract data pairs $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$ as described before, see Figure 2. The whole data set contains $5623$ observation pairs in $19$ different regions, but we only show $2184$ of these pairs. The region in the foreground (black dots, green lines) corresponds to the region with red boundary in Figure 1. It contains $2656$ observations $(\boldsymbol{X}_{\!i},\boldsymbol{V}_{\!\!i})$, and we show a subset of $200$ pairs.
The task is to smooth these noisy data and to determine for various points $\boldsymbol{x}_{o}\in\mathbb{S}^{2}$ whether there is a preferred axis direction of ice cracks close to $\boldsymbol{x}_{o}$. The existence and strength of preferred directions allows conclusions about physical properties of the moon, e.g. the link of tidal forces to ice cracks (Rhoden and Hurford, 2013). Translated into a mathematical task, we want to determine for such a point $\boldsymbol{x}_{o}$ a Bingham distribution (Bingham, 1974) on the unit circle $\mathbb{S}^{2}\cap\boldsymbol{x}_{o}^{\perp}$ of the tangent space $\boldsymbol{x}_{o}^{\perp}$ of the unit sphere at $\boldsymbol{x}_{o}$.
@@
Two general references about directional data, including regression methods, are the monograph of Mardia and Jupp (2000) and the review paper of Pewsey and García-Portugués (2021). These references describe and cite numerous approaches to regression with directional response. Some of these approaches involve transformations of the directional response into points in some Euclidean space, but the choice of this transformation is somewhat ad hoc. Let us restrict our attention to models in which $\boldsymbol{Y}_{\!\!i}\in\mathbb{S}^{d-1}$ has a von Mises–Fisher distribution conditional on $\boldsymbol{X}_{\!i}$. Many authors characterize a von Mises–Fisher distribution via a direction parameter $\boldsymbol{\beta}\in\mathbb{S}^{d-1}$ and a concentration parameter $\kappa\geq 0$ which is often viewed as a nuisance parameter, whereas $\boldsymbol{\beta}$ is the parameter of interest. We propose to replace such a parameter pair $(\boldsymbol{\beta},\kappa)$ with the vector $\boldsymbol{z}=\kappa\boldsymbol{\beta}\in\mathbb{R}^{d}$. Since $\kappa=\|\boldsymbol{z}\|$ and $\boldsymbol{\beta}=\kappa^{-1}\boldsymbol{z}$, this description is equivalent, and dealing with parameters in a Euclidean space is more convenient, particularly in generalized linear models.
The remainder of this paper is structured as follows. Section 2 starts with a short recap of von Mises–Fisher (vMF) distributions and Bingham (Bh) distributions on the unit sphere $\mathbb{S}^{d-1}$ of $\mathbb{R}^{d}$. Then we describe a very useful connection between these two families for dimension $d=2$. Section 3 describes regression methods for data with directional response vectors. As a parametric approach, we propose generalized linear models (GLMs). For a thorough introduction and overview of this topic we refer to McCullagh and Nelder (1989). Thereafter we describe a modification of this parametric approach to local parametric modelling in the spirit of Fan et al. (1998). The latter method is illustrated with simulated data. In Section 4 we explain how to modify the nonparametric methods of Section 3 for smoothing axial data on a sphere, illustrating this method with the specific data mentioned in the beginning. A key tool are stereographic projections of the covariate vectors $\boldsymbol{X}_{\!i}\in\mathbb{S}^{2}$ onto $\mathbb{R}^{2}$ and suitable transformations of the accompanying axes $\boldsymbol{V}_{\!\!i}\in\mathbb{S}^{2}\cap\boldsymbol{X}_{\!i}^{\perp}$ into points on the unit circle $\mathbb{S}^{1}$.
An appendix contains additional information and proofs. In particular, Appendix A provides details about the numerical computation of the mean vector and covariance matrix of a von Mises–Fisher distribution. This involves exact formulae as well as approximations for parameter vectors with large norm and is potentially of independent interest.
@@ 2.1 Von Mises–Fisher distributions
The family $(\mathrm{vMF}(\boldsymbol{z}))_{\boldsymbol{z}\in\mathbb{R}^{d}}$ is a natural exponential family (Barndorff-Nielsen, 2014), and it is well-known that for a random vector $\boldsymbol{Y}\sim\mathrm{vMF}(\boldsymbol{z})$, its mean vector and covariance matrix are given by the gradient and Hessian matrix of $\gamma$,
@@
Other well-known facts are that the mapping $\boldsymbol{\mu}:\mathbb{R}^{d}\to\{\boldsymbol{x}\in\mathbb{R}^{d}:\|\boldsymbol{x}\|<1\}$ is a diffeomorphism. Precisely,
$\boldsymbol{\mu}(t\boldsymbol{v})\ =\ \tilde{\gamma}_{d}^{\prime}(t)\boldsymbol{v}\quad\text{for}\ \boldsymbol{v}\in\mathbb{S}^{d-1},t\in\mathbb{R},$ (1)
where $\tilde{\gamma}_{d}:\mathbb{R}\to\mathbb{R}$ is an even, strictly convex and analytic function such that $\tilde{\gamma}_{d}(0)=0$, and $\tilde{\gamma}_{d}^{\prime}:\mathbb{R}\to(-1,1)$ is bijective, see Appendix A. By means of a Lagrange argument one can show that the density $f_{\boldsymbol{z}}$ maximizes differential Shannon entropy
@@ 2.3 The special case of $d=2$
If we identify any vector in $\mathbb{R}^{2}$ with a complex number, then $\boldsymbol{y}(\boldsymbol{v})=\boldsymbol{v}^{2}$. In particular, this shows immediately that $\|\boldsymbol{y}(\boldsymbol{v})\|=1$. Hence there is a one-to-one correspondence between axes in $\mathbb{R}^{2}$ and the unit circle $\mathbb{S}^{1}$, and specifying the expectation $\int\boldsymbol{v}\boldsymbol{v}^{\top}\,Q(\mathrm{d}\boldsymbol{v})$ for a distribution $Q$ on $\mathbb{S}^{1}$ is equivalent to specifying the expectation $\int\boldsymbol{y}(\boldsymbol{v})\,Q(\mathrm{d}\boldsymbol{v})$. Note further that any symmetric matrix $\boldsymbol{W}\in\mathbb{R}^{2\times 2}$ with trace 00 may be written as
$\boldsymbol{W}\ =\ \begin{bmatrix}z_{1}&z_{2}\\ z_{2}&-z_{1}\end{bmatrix}$ (3)
@@
Furthermore, if $\boldsymbol{V}\sim M$, then $\boldsymbol{y}(\boldsymbol{V})\sim M$ too, and these findings imply the following very useful fact.
Proposition 1.
With $\boldsymbol{W}$ and $\boldsymbol{z}$ as in (3), a random vector $\boldsymbol{V}$ follows $\mathrm{Bh}(\boldsymbol{W})$ if and only if $\boldsymbol{y}(\boldsymbol{V})$ has distribution $\mathrm{vMF}(\boldsymbol{z})$.
@@
$\tilde{f}_{\boldsymbol{w}}(\theta)\ :=\ e^{\kappa\cos(2(\theta-\beta))-\tilde{\gamma}_{2}(\kappa)}$ (4)
@@
$\tilde{\gamma}_{2}(\kappa)\ :=\ \log\Bigl{(}\frac{1}{2\pi}\int_{0}^{2\pi}e^{\kappa\cos(2(\theta-\beta))}\,\mathrm{d}\theta\Bigr{)}\ =\ \log\Bigl{(}\frac{1}{\pi}\int_{0}^{\pi}e^{\kappa\cos(t)}\,\mathrm{d}t\Bigr{)}.$
@@
$\boldsymbol{w}\ :=\ \kappa[\cos(\beta),\sin(\beta)]^{\top}$ (5)
@@
Figure 3 depicts the distribution $\mathrm{Bh}(\boldsymbol{w})$ for two different vectors $\boldsymbol{w}=\kappa[\cos(\beta),\sin(\beta)]^{\top}$. One possible representation is a graph of the angular density $\tilde{f}$ in (4). Another possibility is to draw an “axial histogram”, that is, the region surrounded by the curve $\theta\mapsto\tilde{f}(\theta)^{1/2}[\cos(\theta),\sin(\theta)]^{\top}$, leading to a symmetric but non-circular “pie” (unless $\kappa=0$). The area of any “slice” of this pie with boundary angles $0\leq\theta_{1}<\theta_{2}\leq 2\pi$ equals $2\pi$ times the probability that $\tilde{V}\in[\theta_{1},\theta_{2})$.
@@
The next proposition provides some additional formulae which will be used later and follow from Proposition 1, basic properties of von Mises–Fisher distributions and elementary calculations.
Proposition 2.
Let $\boldsymbol{V}\sim\mathrm{Bh}(\boldsymbol{w})$ with $\boldsymbol{w}$ as in (5). Then,
[…]
with $\boldsymbol{u}:=[\cos(\beta),\sin(\beta)]^{\top}$, the preferred axis direction of $\mathrm{Bh}(\boldsymbol{w})$. Moreover, with $\|\cdot\|_{F}$ denoting Frobenius norm,
$\operatorname{\mathbb{E}}\bigl{(}\|\boldsymbol{V}\boldsymbol{V}^{\top}-\operatorname{\mathbb{E}}(\boldsymbol{V}\boldsymbol{V}^{\top})\|_{F}^{2}\bigr{)}\ \ =\ \frac{1-\tilde{\gamma}_{2}^{\prime}(\kappa)^{2}}{2}.$
In view of this proposition, in the context of regression with Bingham-distributed response vectors, we visualize $\mathrm{Bh}(\boldsymbol{w})$ simply by the vector pair ±$\pm\tilde{\gamma}_{2}^{\prime}(\kappa)\boldsymbol{u}$.
@@ 3 Regression methods for directional data
3.1 Generalized linear models (GLMs)
@@
In our specific settings, any $\boldsymbol{f}\in\mathcal{F}$ can be written as $\boldsymbol{f}=(f_{k})_{k=1}^{d}$ with functions $f_{k}$ in a given finite-dimensional space $\mathcal{F}^{o}$ of functions $f^{o}:\mathcal{X}\to\mathbb{R}$. For this particular case, Appendix B provides technical details about the first and second derivatives of the function $\ell(\cdot)$ after a suitable parametrization of $\mathcal{F}_{o}$. These formulae enable us to minimize $\ell(\boldsymbol{f})$ over all $\boldsymbol{f}\in\mathcal{F}$ via a Newton-Raphson procedure.